If a thin film has a thickness that is

A. 1/4 of a wavelength, constructive interference could occur.
B. 1/4 of a wavelength, destructive interference could occur.
C. 1/2 of a wavelength, constructive interference could occur.
D. 1/2 of a wavelength, destructive interference could occur.
Answer #1

1/4 of a wavelength, constructive interference could occur.

Option (a) is correct.

Because ,

Constructive interference occurs here when

2{t}_{\text{c}}=\frac{{\lambda }_{n}}{2},\phantom{\rule{0.25em}{0ex}}\frac{{3\lambda }_{n}}{2},\phantom{\rule{0.25em}{0ex}}\frac{{5\lambda }_{n}}{2}\text{,}\dots .

The smallest constructive thickness {t}_{\text{c}} thus is

t= \lambda /4

This is the condition for constructive interference....